Of 100 staff members of a school, 48 drink coffee, 25 drink both tea & coffee& everyone drinks either tea or coffee, how many drink tea?
Answer by jim_thompson5910(35256)

"48 drink coffee, 25 drink both tea & coffee"

so 48-25 = 23 people just drink coffee

x = # of people who just drink tea

(# of people who drink just coffee)+(# of people who drink just tea)+(# of people who drink both) = total

23+x+25 = 100

x+48 = 100

x = 100 - 48

x = 52

There are 52 people who drink tea only.

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------